40 equal end-of-year deposits are made into a savings account that pays 1% interest. Compute the amount of each deposit that will permit withdrawals of $20,000 at the ends of the last 5 years, leaving the account empty.
Answer #1

Future value of withdrawals after 5 years ($) = 20,000 x F/A(1%, 5) = 20,000 x 5.101** = 102,020

Value of equal year-end deposits ($) = 102,020 / F/A(1%, 40) = 102,020 / 48.8864** = 2,086.88

**From F/A factor table